Troy Engines Ltd. manufactures a variety of engines for use in heavy equipment. The company has always produced all of the necessary parts for its engines, including all of the carburetors. An outside supplier has offered to produce and sell one type of carburetor to Troy Engines Ltd. for a cost of $75.0 per unit. To evaluate this offer, Troy Engines Ltd. has gathered the following information relating to its own cost of producing the carburetor internally:

  

Per
Unit
30,000 Units
per Year
  Direct materials $ 19 $ 570,000
  Direct labour 20 600,000
  Variable manufacturing overhead 13 390,000
  Fixed manufacturing overhead, traceable 21.0 * 630,000
  Fixed manufacturing overhead, allocated 19 570,000
  Total cost $ 92.0 $ 2,760,000
* One-third supervisory salaries; two-thirds depreciation of special equipment (no resale value)
1-a.

Compute the total differential cost per unit for producing and buying the product.

make buy
total differential cost(per unit)

2-a.Suppose that if the carburetors were purchased, Troy Engines Ltd. could use the freed capacity to launch a new product. The segment margin of the new product would be $500,000 per year. Compute the total differential cost in producing and buying the product when the segment margin is foregone on a potential new product line.

make buy
total differential cost
Answer #1

1a) Differential Cost per unit

Make Buy
Direct material 19
Direct labor 20
Variable manufacturing overhead 13
Fixed manufacturing overhead, traceable (21/3) 7
Purchase cost 75
Total differential cost per unit 59 75

2a) Differential Cost

Make Buy
Direct material 19*30000 = 570000
Direct labor 20*30000 = 600000
Variable manufacturing overhead 13*30000 = 390000
Fixed manufacturing overhead, traceable (21/3) 7*30000 = 210000
Opportunity Cost 500000
Purchase cost 75*30000 =
Total differential cost per unit 2270000 2250000